Given f(x)=3x - 1, find
a. f(2)
b. f(-1)
c. f(0)
d. f(2/3)
e. f(k)
f. f(k + 2)

Answer by checkley77(12844)

Given f(x)=3x - 1
f(2)=3*2-1
f92)=6-1
f(2)=5 ans.\
f(-1)=3*-1-1
f(-1)=-3-1
f(-1)=-4 ans.
f(0)=3*0-1
f(0)=0-1
f(0)=-1 ans.
I'll leave the rest for you to work on.